# Temporal Triplane Transformers as Occupancy World Models

### v1

@@ 4.3 Visualizations
Comparisons with SOTA. Figure 4 illustrates the evolution of the scene over the next 3 seconds, along with the corresponding motion predictions for two different scenarios. (i) In the first case, it is clear that OccWorld-O incorrectly labels the truck as a trailer initially, resulting in a continuous accumulation of errors. In contrast, T3Former effectively learns the motion dynamics of all objects in the scene, consistently tracking each vehicle over a long horizon, while providing clearer object boundaries. (ii) In the second case, as the prediction horizon extends, OccWorld-O’s road boundary gradually fades, losing its clarity, whereas our method maintains the boundary shape of the road, including the trees. Additionally, OccWorld-O misidentifies the motorcycle as a truck at both 2.5s and 3s, while T3Former-O ensures that the semantics of small objects remain accurate over time. (iii) Under both scenarios, T3Former-O enables precise long-term future predictions, allowing our motion planning to rely solely on the evolving scene dynamics. This leads to more accurate trajectories and better avoidance of potential future hazards.
@@
Figure 6: Predicted locations of dynamic objects for the three last frames in the prediction window.
Figure 7: A failure case at a T-junction scenario.
Comparison of motion prediction. Figure 6 compares occupancy predictions for the three furthest frames in the prediction window, focusing on dynamic objects of varying sizes (cars and pedestrians) from a BEV perspective. Two reference lines are displayed to reflect the errors between the predictions and GT. The results of our method match the GT significantly better than OccWorld-O, which exhibits substantial drift. For instance, the predictions for the pedestrians at the top and the vehicle on the left clearly demonstrate this improvement
Failure case. Figure 7 shows a failure case across three frames, where the road boundary at the end contains sparse occupied voxels, making the prediction challenging. Due to the inherent blurring issue of VAE [56, 57], OccWorld predictions tend to smooth and homogenize the sparse regions. In contrast, the triplane consists of fine-grained geometric details, which causes our predictions to also exhibit sparse boundaries, resulting in some deviation in the trajectory.
@@ 4.4 Ablation study
To validate the effectiveness of the key designs in T3Former, we conduct ablation studies by evaluating the averaged mIoU and averaged L2 across future outcomes, as shown in Table 3. M0 represents our complete version. For fairness, all ablation models are derived by replacing specific modules in M0, as outlined below: M1 replaces the input of T3Former with the VQ-VAE token from OccWorld [3], which results in a significant performance drop. M2 uses a single-scale Transformer to learn temporal dynamics based on triplane representations, which fails to capture the motion patterns of multi-scale objects, leading to poor performance. M3 leverages multi-scale Transformers for parallel forecasting, as shown in Figure 1(a), instead of autoregression, which leads to error accumulation. M4 directly predicts the entire future triplane, demonstrating that predicting changes in the triplane is more efficient.
Idx. Models Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$
M0 T3Former-O 34.70 0.63
M1 w/o Triplane 25.76 0.96
M2 w/o Multi-scale mot. 31.92 0.87
M3 w/o Autoregression 32.44 0.79
M4 w/o Changing triplane 33.81 0.71
Table 3: Ablation study of the components in T3Former.
4.5 Efficiency analysis
The predefined latent space shapes of OccWorld [3] and OccLLM [5] are 128$\times$50$\times$50 and 64$\times$50$\times$50, respectively. In contrast, T3Former only requires a compact triplane representation that is in spirit similar to orthogonal decomposition [33, 32]. Specifically, the $xy$ plane has a shape of 8$\times$100$\times$100, while the $xz$ and $yz$ planes are represented with shapes of 8$\times$100$\times$16. This results in the smallest latent shape and parameter count for $\Phi_{\rm enc,dec}$, while achieving superior reconstruction accuracy, as shown in Table 4.
Models Latent shape $\downarrow$ Params (M) $\downarrow$ Reconstruction
mIoU $\uparrow$ IoU $\uparrow$
OccWorld [3] 320,000 14.17 60.50 59.07
OccLLM [5] 160,000 2.30 71.08 62.74
T3Former 105,600 0.69 85.50 92.07
Table 4: Comparisons of latent representations within OWMs.
Based on the triplane representation learned by $\Phi_{\rm enc}$, our $\Phi_{\rm fut}$ only needs to learn the changes in the triplane, while $\Phi_{\rm act}$ predicts the motion trajectory based on these changes. This results in our lightweight design, with a total parameter count of 55M, compared to OccWorld’s 72M, achieving a 23% reduction. As shown in Figure 1(c), T3Former achieves 26 FPS for predicting each future frame on an RTX 4090 GPU, whereas OccWorld delivers only 18 FPS.
@@ 5 Conclusion
This paper proposes a new 4D occupancy world model, T3Former, which features three key designs: (i) pre-training triplane representations for OWM, (ii) capturing comprehensive motion dynamics using multi-scale Transformers, and (iii) predicting triplane changes for occupancy forecasting and motion planning. T3Former achieves SOTA performance in 4D occupancy forecasting and motion planning, with real-time capability. Our future work will dive deeper into the generation of driving views, as shown in Figure 5, focusing on the synergy between prompt constraints, image texture, and T3Former.

### v2

@@ 4.3 Visualizations
Comparisons with SOTA.
Figure 6 illustrates the evolution of the scene over the next 3 seconds, along with the corresponding motion predictions for two different scenarios. (i) In the first case, it is clear that OccWorld-O incorrectly labels the truck as a trailer initially, resulting in a continuous accumulation of errors. In contrast, T3Former effectively learns the motion dynamics of all objects in the scene, consistently tracking each vehicle over a long horizon, while providing clearer object boundaries. (ii) In the second
Figure 8: Predicted locations of dynamic objects for the three last frames.
case, as the prediction horizon extends, OccWorld-O’s road boundary gradually fades, losing its clarity, whereas our method maintains the boundary shape of the road, including the trees. Additionally, OccWorld-O misidentifies the motorcycle as a truck at both 2.5s and 3s, while T3Former-O ensures that the semantics of small objects remain accurate over time. (iii) Under both scenarios, T3Former-O enables precise long-term future predictions, allowing our motion planning to rely solely on the evolving scene dynamics. This leads to more accurate trajectories and better avoidance of potential future hazards.
@@
Comparison of motion prediction. Figure 8 compares occupancy predictions for the three furthest frames in the prediction window, focusing on dynamic objects of varying sizes (cars and pedestrians) from a BEV perspective. Two reference lines are displayed to reflect
Figure 9: Limitations in sparse boundary regions.
Table 3: Ablation study of T3Former.
Idx. Models Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$ FPS $\uparrow$
M0 T3Former-O 36.09 1.00 26
M1 w/o pretraining 33.96 1.17 26
M2 w/o triplane 28.97 1.10 21
M3 w/o triplane changes 30.51 1.04 27
M4 w/o multi-scale mot. 31.92 1.12 36
M5 w/o autoregression 34.65 1.06 34
Table 4: Effect of different hyperparameters.
Idx. Setting Avg. mIoU $\uparrow$ Avg. L2 $\downarrow$ FPS $\uparrow$
H1.1 8, (8, 50, 50) 26.78 1.33 37
H1.2 8, (16, 50, 50) 28.91 1.23 33
H1.3 8, (8, 100, 100) 31.42 1.17 30
H1.4 8, (16, 100, 100) 36.09 1.00 26
H1.5 16, (16, 100, 100) 36.20 1.09 23
H1.6 32, (16, 100, 100) 35.71 1.11 18
H2.1 $V=3$, (4, 16) 35.91 1.08 28
H2.2 $V=5$, (4, 16) 36.09 1.00 26
H2.3 $V=5$, (4, 32) 36.14 1.04 21
H2.4 $V=5$, (8, 16) 36.18 1.02 23
H2.5 $V=7$, (4, 16) 35.65 0.98 20
the errors between the predictions and GT. The results of our method match the GT significantly better than OccWorld-O, which exhibits substantial drift. For instance, the predictions for the pedestrians at the top and the vehicle on the left clearly demonstrate this improvement
Limitation. As shown in Figure 9, T3Former produces sparse predictions in boundary regions. This is primarily due to occlusions in the original occupancy data, coupled with the use of triplanes that employ the orthogonal decomposition strategy [25, 17] to encode fine-grained geometric details. While this approach is effective for detailed representation, it complicates the task of correlating surrounding context and filling in the sparse areas during forecasting. In contrast, OccWorld tends to smooth and homogenize these sparse regions, a result of the inherent blurring issue in VAE-based models [5, 12]. In future work, we aim to improve content generation in these sparsely observed regions.
@@ 4.4 Ablation study
Table 3 validates the effectiveness of the key designs in T3Former. M0 represents our complete version. For fairness, all ablation models are derived by replacing specific modules in M0, as outlined below: M1 directly learns triplanes and downstream tasks (Eqs. (9) and (11)) in an end-to-end manner, without pre-training, leading to reduced encoder efficiency and performance degradation. M2 replaces T3Former’ input with the VQ-VAE token from OccWorld [58], causing some geometric information loss. M3 predicts the entire future triplane instead of $\Delta$ changes, showing that change prediction is more efficient. M4 uses a single-scale Transformer to model temporal dynamics, failing to capture multi-scale object motion and causing error accumulation. M5 predicts all future outcomes in parallel, resulting in poor performance due to the inability to correct autoregressive errors.
Table 4 analyzes the effect of different hyperparameters. denotes our trade-off version. The H1 series illustrates the impact of different latent shapes of triplanes. The setting refers to the number of channels, as well as the (height, width, and length). It is clear that increasing the spatial shape significantly enhances future prediction performance, while larger channel numbers provide little benefit and may even reduce performance. This is due to the introduction of feature redundancy, which complicates self-attention learning during forecasting. In the H2 series, the setting defines the number of scales, Transformer layers, and self-attention heads in T3Former. Our results show that even a lightweight Transformer effectively captures key patterns, while increasing model size brings little benefit but adds computational overhead. Consequently, our design uses only 55M parameters – 23% fewer than OccWorld’s 72M – yet achieves superior efficiency. As illustrated in Figure 2, T3Former runs at 26 FPS on an RTX 4090 GPU, outperforming OccWorld’s 18 FPS.
@@ 5 Conclusion
This paper presents T3Former, a new 4D occupancy world model with three key designs: (i) pre-trained triplane representations for OWM, (ii) multi-scale Transformers for capturing motion dynamics, and (iii) triplane change prediction for occupancy forecasting and motion planning. T3Former achieves SOTA performance in 4D occupancy forecasting and motion planning with real-time capability. Our future work will dive deeper into driving view generation (see Figure 8), focusing on the synergy between prompt constraints, image texture, and temporal consistency in occupancy data.